# Subtleties in the pseudomodes formalism

### 04 Mar 2026 v2

@@ Acknowledgments
This research is primarily supported by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES) under Award No. DE-SC0025516. L.P.B. acknowledges Research Ireland for support through the Frontiers for the Future project. Optimization was performed with the Optim.jl Julia package Mogensen and Riseth (2018). No generative AI tools were used in the writing of this article.
@@ Appendix A Derivation of the Effective Spectral Density
where ${|{\mathbf{1}_{l}}\rangle}=a_{l}{|{\mathbf{0}}\rangle}$. Now define a vector $v_{i}$ such that its components are $v_{i,j}(t)={\langle{\mathbf{0}}|}a_{i}(t){|{\mathbf{1}_{j}}\rangle}$ and consider:
$\frac{dv_{i,j}}{dt}$ $={\langle{0}|}\dot{a}_{i}(t){|{1_{j}}\rangle}={\langle{0}|}\mathcal{L}^{\prime}({a}_{i}(t)){|{1_{j}}\rangle}\,$, (126)
@@
${\langle{\mathbf{0}}|}[H_{E},a_{i}(t)]{|{\mathbf{1}_{j}}\rangle}$ $=0+\sum_{m=1}^{n_{\alpha}}\Lambda_{mj}{\langle{\mathbf{0}}|}a_{i}(t){|{\mathbf{1}_{m}}\rangle}\,$, (127)
@@ Appendix B Memory Kernel Fitting Methods
for complex parameters $\kappa_{k},\lambda_{k}$, where we want $n$ as small as possible to get a good fit. In this section we describe two algorithms from discrete signal processing, Prony’s method and ESPRIT, that have been used for this fitting problem.
@@ B.1 Prony’s Method
Prony’s method de Prony (1795); Wilson et al. (2019); Almunif et al. (2020); Chen et al. (2022) consists of the following steps:
@@ Appendix C Non-Diagonalizable Configuration as the Limit of a Diagonalizable One
Consider the 2x2 non-diagonalizable block from section II.3 where we set $\varepsilon=\eta=0$, and apply a small perturbation $\epsilon$ to the coupling between the two modes:
@@
Appendix D 3x3 non-diagonalizable Configuration
In general, constructing an $n\times n$ non-diagonalizable matrix that is the sum of an anti-Hermitian matrix a and real diagonal matrix, as $W$ is, is a nontrivial task. However, we can use the 2x2 non-diagonalizable block to construct a 3x3 non-diagonalizable block for 3 pseudomodes arranged in a chain. We begin with the 2x2 non-diagonalizable block and expand it to 3x3, leaving the two additional off-diagonal elements as unknown:
@@
and so we find a candidate 3x3 non-diagonalizable $W$ as:
@@
If we then proceed as in the 2x2 case and allow only the first pseudomode to couple to the system and the last pseudomode to couple to the residual bath (so $\eta=0$), we can then compute the resulting effective memory kernel
@@ Appendix E Inversion Problem for 2 Pseudomodes
In this appendix we explore the simple case of two pseudomodes. While a solution for the pseudomode parameters in this case has been previously shown Pleasance et al. (2020), we are able to show using the methods of section III that, for the 2-mode case, it is always possible to choose pseudomode parameters such that $\Gamma>0$ for effective spectral density that is nonnegative.
@@
Since the matrices here are only 2x2, we can then construct $S^{\dagger}S$ directly from (45) as:
@@
$\alpha^{2}r^{2}\gamma^{2}>4\beta^{2}\varepsilon^{2}\,.$ (171)
@@ Appendix F Fit Parameters in Figures 2 and 4
Table 2: Fit parameters for the diagonal fit used in Fig. 2(a,b), as they appear in (20).
@@
Table 4: Fit parameters for the diagonalizable fit used in Fig. 2(a,c), as they appear in (27).
$\kappa_{k}$ $\varepsilon_{k}$ $\gamma_{k}$
0.67900 + 0.37224ii -0.34457 -1.23329
0.67900 - 0.37224ii 0.34457 -1.23329
0.10729 - 0.32214ii 0.79313 -0.62406
0.10729 + 0.32214ii -0.79313 -0.62406
-0.00089 - 0.07379ii 0.95949 -0.15409
-0.00089 + 0.07379ii -0.95949 -0.15409
$\kappa_{k}$ $\varepsilon_{k}$ $\gamma_{k}$
0.0369 -0.706 0.295
0.0369 0.706 0.295
0.0699 -0.377 0.423
0.0699 0.377 0.423
[…]

### 22 May 2026 v3

@@ Acknowledgments
This research is primarily supported by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES) under Award No. DE-SC0025516. L.P.B. acknowledges Research Ireland for support through the Frontiers for the Future project. Optimization was performed with the Optim.jl Julia package Mogensen and Riseth (2018). Numerical data used in this article are provided in Appendix F. No generative AI tools were used in the writing of this article.
@@ Appendix A Derivation of the Effective Spectral Density
where ${|{\mathbf{1}_{l}}\rangle}=a_{l}^{\dagger}{|{\mathbf{0}}\rangle}$. Now define a vector $v_{i}$ such that its components are $v_{i,j}(t)={\langle{\mathbf{0}}|}a_{i}(t){|{\mathbf{1}_{j}}\rangle}$ and consider:
$\frac{dv_{i,j}}{dt}$ $={\langle{0}|}\dot{a}_{i}(t){|{1_{j}}\rangle}={\langle{0}|}\mathcal{L}^{\prime}{a}_{i}(t){|{1_{j}}\rangle}\,$, (125)
@@
${\langle{\mathbf{0}}|}[H,a_{i}(t)]{|{\mathbf{1}_{j}}\rangle}$ $=0+\sum_{m=1}^{n_{\alpha}}\Lambda_{mj}{\langle{\mathbf{0}}|}a_{i}(t){|{\mathbf{1}_{m}}\rangle}\,$, (126)
@@ Appendix B Memory Kernel Fitting Methods
for complex parameters $\kappa_{k},\lambda_{k}$, where we want $n$ as small as possible to get a good fit. In this appendix we describe two algorithms from discrete signal processing, Prony’s method and ESPRIT, that have been used for this fitting problem.
@@ B.1 Prony’s Method
Prony’s method Prony et al. (1795); Wilson et al. (2019); Almunif et al. (2020); Chen et al. (2022) consists of the following steps:
@@ Appendix C Non-Diagonalizable Configuration as the Limit of a Diagonalizable One
Consider the $2\times 2$ non-diagonalizable block from section II.3 where we set $\varepsilon=\eta=0$, and apply a small perturbation $\epsilon$ to the coupling between the two modes:
@@
Appendix D $\mathbf{3\times 3}$ non-diagonalizable Configuration
In general, constructing an $n\times n$ non-diagonalizable matrix that is the sum of an anti-Hermitian matrix a and real diagonal matrix, as $W$ is, is a nontrivial task. However, we can use the $2\times 2$ non-diagonalizable block to construct a $3\times 3$ non-diagonalizable block for 3 pseudomodes arranged in a chain. We begin with the $2\times 2$ non-diagonalizable block and expand it to $3\times 3$, leaving the two additional off-diagonal elements as unknown:
@@ Appendix D $\mathbf{3\times 3}$ non-diagonalizable Configuration
and so we find a candidate $3\times 3$ non-diagonalizable $W$ as:
@@
If we then proceed as in the $2\times 2$ case and allow only the first pseudomode to couple to the system and the last pseudomode to couple to the residual bath (so $\eta=0$), we can then compute the resulting effective memory kernel
@@ Appendix E Inversion Problem for 2 Pseudomodes
In this appendix we explore the simple case of two pseudomodes. While a solution for the pseudomode parameters in this case has been previously shown Pleasance et al. (2020), we are able to show using the methods of Sec. III that, for the 2-mode case, it is always possible to choose pseudomode parameters such that $\Gamma>0$ for effective spectral density that is nonnegative.
@@
Since the matrices here are only $2\times 2$, we can then construct $S^{\dagger}S$ directly from (44) as:
@@
$r^{2}\alpha^{2}\gamma^{2}>4\beta^{2}\varepsilon^{2}\,.$ (170)
@@ Appendix F Fit Parameters in Figures 2 and 4
All numerical data in this article are shown, or can be reproduced from, the data contained in the following tables.
Table 2: Fit parameters for the diagonal fit used in Figs. 2(a) and 2(b), as they appear in (19).
@@
Table 3: Fit parameters for the diagonalizable fit used in Figs. 2(a) and 2(c), as they appear in (26).
$\kappa_{k}$ $\varepsilon_{k}$ $\gamma_{k}$
0.10807 + 0.05924ii 0.34457 1.23329
0.10807 - 0.05924ii -0.34457 1.23329
0.01708 - 0.05127ii -0.79313 0.62406
0.01708 + 0.05127ii 0.79313 0.62406
0.00014 - 0.01174ii -0.95949 0.15409
0.00014 + 0.01174ii 0.95949 0.15409
$\kappa_{k}$ $\varepsilon_{k}$ $\gamma_{k}$
0.0369 -0.706 0.295
0.0369 0.706 0.295
0.0699 -0.377 0.423
0.0699 0.377 0.423
[…]